# HoloVIC: Large-scale Dataset and Benchmark for Multi-Sensor Holographic Intersection and Vehicle-Infrastructure Cooperative

### v2

@@ 4.2 Multi-Sensor Perception
where $GT_{s}$ is the ground truth in $s$-th Sensor, $S$ is the set of Sensors in Intersection. We refer to 3D Detection metrics [9] and adopt mAP and mAOS as the evaluation metrics for multi-sensor 3D detection. In addition, we utilized metrics of Multi-target Multi-camera Tracking (MTMCT) [19] and selected MOTA and IDF1 as the metrics for multi-sensor 3D Tracking. All the definitions of metrics are consistent with the relevant descriptions in Eq.33-39.
@@ Appendix A Coordinates Transformation
Table 7: The Number and Ratio of 3D Boxes (The total from all sensors) and BEV Boxes (Global Scenes) per category.
Local/Global Vechile Cyclist Pedestrian Large Vehicle Total
Local (3D Boxes) 7.45M 3.26M 734k 27.5k 11.47M
64.92% 28.44% 6.4% 0.24% 100%
Global (BEV Boxes) 2.39M 941k 256k 13.7k 3.60M
20.87% 8.2% 2.23% 0.12% 31.42%
@@
Appendix B More Details of HoloVIC
We analyze the number and ratio of 3D boxes for all sensors and BEV scenes as shown in Tab.7. Due to some large vehicles such as buses, trucks, engineering vehicles, etc. often appear in the scene, we specially divided the vehicles with length greater than 7 meters into a separate category ”Large Vehicle”. The ratio of vehicles, cyclist, pedestrian and large vehicle are 64.92%, 28.44%, 6.4% and 0.24%, respectively, and the vehicles are still the most targets in the traffic scene. Since some targets are captured by 2-6 sensors, the number of boxes in BEV scenes is less than the total number of boxes from sensors, which the BEV boxes are accounted for 31.42% of the local boxes from sensors.
The distribution of all intersections in the HoloVIC Dataset in HD-Map is shown in Fig.A2. The red dashed box identifies the corresponding intersection number for each intersection (Int-1/VIC-1)-(Int-5/VIC-5), which share the same Global Coordinate System. The illustration of intersections with different sensor layouts (Type A-D) are shown in Fig.A3-A5. The illustration of Vehicle-Infrastructure Cooperative is illustrated in Fig.A7.

### v3

@@ 4.2 Multi-Sensor Perception
where $GT_{s}$ is the ground truth in $s$-th Sensor, $S$ is the set of Sensors in Intersection. We refer to 3D Detection metrics [9] and adopt mAP and mAOS as the evaluation metrics for multi-sensor 3D detection. In addition, we utilized metrics of Multi-target Multi-camera Tracking (MTMCT) [19] and selected MOTA and IDF1 as the metrics for multi-sensor 3D Tracking. All the definitions of metrics are consistent with the relevant descriptions in Eq.4-10.
@@ Appendix A Coordinates Transformation
Appendix B Annotation Process
B.1 Device Time Synchronization
All devices at each intersection are connected to a switch, and their time synchronization is achieved through NTP (Network Time Protocol), which ensure the time error is less than 5ms. When collecting data, the cameras and fisheyes capture at a frequency of 25Hz, while the LiDAR operates at 10Hz.
To construct the dataset, we establish timestamp anchors at a frequency of 10Hz along the timeline. We then select the nearest frame from each device around each anchor and package them into frame batches. Each batch contains synchronized data from all devices at that specific time, and is assigned a corresponding frame index.
B.2 Calibration
Once the sensors are deployed at each intersection and time synchronization is complete, we need to calibrate all sensors. For cameras and fisheyes, calibration involves determining the distortion, intrinsic, and extrinsic parameters. For LiDARs, only the extrinsic parameters need to be calibrated. These extrinsic parameters establish the transformation relationship between the intersection/vehicle coordinate system and the device coordinate system, as explained in Sec. A.
To make it easier for researchers to use our dataset, all images captured by the cameras and fisheyes are undistorted. The coordinate transformation is based on undistorted pixel coordinates and the intersection/vehicle coordinate system. Furthermore, the coordinate transformation between devices can be achieved by linking their extrinsic parameters to the intersection/vehicle coordinate system.
B.3 Global Annotation
We merge all the point clouds within each frame batch. Using Eq.(3) from the supplementary material, we project each individual point cloud onto the intersection coordinate system. The two sets of point clouds are concatenated together. And then, we annotate the 3D boxes for the targets that appear in the concatenated point cloud scene. This annotation process involves determining the positions, orientations, and categories of the 3D boxes.
Afterwards, the annotated 3D boxes are projected onto the images using the corresponding extrinsic parameters for each camera and fisheye. Annotators are then tasked with performing supplementary annotations for each camera, especially in cases where the target’s point cloud is occluded or extends beyond the range captured by the LiDAR, resulting in missed annotations. The annotated boxes are subsequently projected onto the intersection/vehicle coordinate system. However, it is important to note that due to calibration errors, there may be a minuscule number of boxes in the dataset that have inaccurate projections onto the intersection position.
Annotators associate a global ID with the annotated 3D boxes in the timeline. Subsequently, all the 3D boxes are reprojected onto all the devices. Annotators then determine the visibility of each box, indicating which devices can see the box. For example, if a box is only visible in Lidar-2, Camera-1, Camera-3, and Fisheye-1, then the visibility information for that box will have ”True” (”Visible”) for those devices and ”False” (”Invisible”) for the rest. There are several situations where a box may be marked as ”Invisible”: If the global 3D box is outside the field of view of a device; If the object is occluded by other objects that exceeds more than 80%; If the object appears too small in the image (far from the device).
Considering that many researchers often use trajectory inpainting based on temporal to handle occlusions in tracking tasks, it is worth noting that even though an object may be occluded and not visible in a specific frame, the inpainted boxes can still accurately outline its correct position. Therefore, during evaluation, all of the boxes are labels as ”Invisible” are not counted as false positives or false negatives for calculating mAP, MOTA, IDF1, etc., regardless of whether the prediction boxes provided or not.
Both the vehicle-side and the road-side undergo the calibration process described above. Afterwards, all the 3D boxes are transformed onto the global coordinate system. The global ID association is then determined based on the Intersection over Union (IOU) between the 3D boxes of the vehicles and the road.
Appendix C Visualization of HoloVIC
We show all of the visualization results involving all of intersections Int-1/VIC-1 to Int-5/VIC-5. The distribution of all intersections in the HoloVIC Dataset in HD-Map is shown in Fig.A3. The red dashed box identifies the corresponding intersection number for each intersection (Int-1/VIC-1)-(Int-5/VIC-5), which share the same Global Coordinate System. The illustration of Vehicle-Infrastructure Cooperative is illustrated in Fig.A2. The illustration of intersections with different sensor layouts (Type A-D) are shown in Fig.A4-A7.